Factor:
42a^2-29ab-5b^2

Answer by checkley75(3666)

42a^2-29ab-5b^2 factor the 42=6*7 factor the 5=1*-5 or 5=-1*5
now multiply the two factors -5*7=-35 & 1*6=6 now add them -35+6=-29
(7a+b)(6a-5b)